Cupcake123456:

Gene paid a deposit on a leased car. The deposit earns 2.8 percent simple annual interest. At the end of the year, the interest that is earned is $22.40. What was the amount of the original deposit? $0.63 $80.00 $62.72 $800.00 help? @thisgirlpretty @shadow @563blackghost @razor @angle

JustSaiyan:

I don't do mathematics. Sorry. I only do History, Biology, and Chemistry.

Cupcake123456:

You know anyone who does ?

Angle:

Hello ^_^ so for simple interests the formula would be: (Deposit)*(number of years)*(interest rate) = (interest earned)

Angle:

here is the information we know: number of years = 1 interest rate = 0.028 interest earned = 22.4

JustSaiyan:

Angle does.

Cupcake123456:

ok

Cupcake123456:

i got the information @angel

Angle:

can you plug it into the formula I gave?

Cupcake123456:

ill try

Ultrilliam:

Note: @angel =/= @Angle

Angle:

show me what you get when you're ready ^_^

Cupcake123456:

i tried already & b

Cupcake123456:

@angle ^

Angle:

hmm well we have (Deposit) * (1) * (0.028) = 22.4 right?

Cupcake123456:

right

Angle:

then Deposit = (22.4) / (0.028) = ?

Cupcake123456:

800?

Angle:

yup!

Cupcake123456:

Thanksss !